At the tennis camp the ratio to boys to girls was 6 to 5.if there were 121 people altogether

Answer by checkley79(3341)

B+G=121
B:G=6:5
B=121-G
(121-G)/G=6/5
6G=5(121-G)
6G=605-5G
6G+5G=605
11G=605
G=605/11
G=55 ANS. FOR THE NUMBER OF GIRLS.
B+5=121
B=121-55
B=66 ANS. FOR THE NUMBER OF BOYS.
PROOF:
66:55=6:5
6:5=6:5